OpenStudy (anonymous):

Which of (3, 5), (4, 6), (5, 7), and (6, 8) are solutions to y = x + 2?

hartnn (hartnn):

the pairs (x,y), check whether they follow, y=x+2 is 5=3+2 ?? 6=4+2 ?? 7=5+2 ?? 8=6+2 ?? whichever answer is YES, are the solution.

OpenStudy (anonymous):

so all of them are right ?

hartnn (hartnn):

yup, all are solutions.

OpenStudy (anonymous):

k , thanks (:

hartnn (hartnn):

welcome :)